# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

### v1

@@ 2 Preliminaries
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable and $rangeF_{*}$ is Ricci flat. Then
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)=\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)\\ &+rg_{M}(\nabla_{CX}\nabla f,CY)+rCX(f)CY(f)-\lambda g_{M}(CX,CY)=0.\end{split}$ (3.22)
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM),$ we get
$\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric(X,Y)+\lambda g_{M}(X,Y)=0.$
Making use of Corollary 3.2, we obtain
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\lambda g_{M}(X,Y)=0.\end{split}$
Since $rangeF_{*}$ is Ricci flat, the above equation reduces to
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)-rCX(f)CY(f)+\lambda g_{M}(CX,CY)=0.\end{split}$
Since $\eta$ is vertical and $(kerF_{*})^{\perp}$ is integrable, we get
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)=\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)\\ &+rg_{M}(\nabla_{CX}\nabla f,CY)+rCX(f)CY(f)-\lambda g_{M}(CX,CY).\end{split}$
This completes the proof. ∎
Theorem 3.7.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable. Then
@@
where $\alpha^{\prime}=\frac{-1}{r}.$
@@
Since $\eta$ is horizontal and $(kerF_{*})^{\perp}$ is integrable, we get
@@
$\frac{1}{2}(L_{F_{*}(\nabla f)}g_{N})(F_{*}JU,F_{*}CX)+\alpha^{\prime}Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=0.$
@@
Let $(M,g_{M},J)$ be a 6-dimensional Kähler manifold, where $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{4}:x_{i}\neq 0,i=1,\ldots,6\},$ with Riemannian metric
$g_{M}=e^{-2x_{3}}dx_{1}^{2}+e^{-2x_{3}}dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}+e^{-2x_{6}}dx_{5}^{2}+dx_{6}^{2},$
@@
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=e_{3},X_{2}=e_{4},X_{3}=e_{5},X_{4}=e_{6}\}.$
Here,
$\{e_{1}=e^{x_{3}}\frac{\partial}{\partial x_{1}},e_{2}=e^{x_{3}}\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}},e_{5}=e^{x_{6}}\frac{\partial}{\partial x_{5}},e_{6}=\frac{\partial}{\partial x_{6}}\}$
@@
$F_{*}(X_{1})=e_{1}^{\prime},$ $F_{*}(X_{2})=-sinx_{4}e_{3}^{\prime}+cosx_{4}e_{4}^{\prime},F_{*}(X_{3})=e_{5}^{\prime},F_{*}(X_{4})=e_{6}^{\prime},$
where $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=\frac{\partial}{\partial y_{6}}\}$ represents the standard orthonormal basis of $T_{F(p)}N.$
Also, it can be easily verified that for every $i=1,2,3,4$
@@
and $J(U_{1})=-X_{4},J(U_{2})=-X_{3},J(X_{1})=-X_{2},J(X_{2})=-X_{1},J(X_{3})=-U_{2},J(X_{4})=-U_{1}.$ Thus $F$ is an anti-invariant Riemannian map.
@@
$\Gamma^{1}_{13}=-1=\Gamma^{1}_{31},\Gamma^{3}_{11}=e^{-2x_{3}},\Gamma^{2}_{23}=-1=\Gamma^{2}_{32},\Gamma^{3}_{22}=e^{-2x_{3}},\Gamma^{5}_{56}=-1=\Gamma^{5}_{65},\Gamma^{6}_{55}=e^{-2x_{6}}$ and rest of them are zero.
@@
$\nabla_{U_{1}}U_{1}=X_{1},\nabla_{U_{1}}X_{1}=-U_{1},\nabla_{U_{2}}U_{2}=X_{1},\nabla_{U_{2}}X_{1}=-U_{2},\nabla_{X_{3}}X_{3}=X_{4},\nabla_{X_{3}}X_{4}=-X_{3},$ rest of them are zero.
Ricci curvatures are obtained as $Ric(X_{2},X_{2})=-2,Ric(X_{3},X_{3})=-2,Ric(X_{4},X_{4})=-2,$ remaining Ricci tensors are zero.
@@
$\begin{split}(L_{Z_{1}}g_{M})(Z_{2},Z_{3})=&\frac{1}{2}(a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+l_{1}l_{2}h_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+l_{1}l_{3}h_{2})\\ &-a_{2}c_{1}a_{3}-b_{3}b_{2}c_{1}-l_{2}l_{3}h_{1},\end{split}$ (3.23)
$Ric(Z_{2},Z_{3})=-2l_{2}l_{3}-2h_{2}h_{3}-2d_{2}d_{3},$ (3.24)
@@
$A_{1}=2(2l_{2}l_{3}+2h_{2}h_{3}+2d_{2}d_{3}+a_{2}a_{3}c_{1}+b_{2}b_{3}c_{1}+l_{2}l_{3}h_{1}),$
$A_{2}=a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+l_{1}l_{2}h_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+l_{1}l_{3}h_{2}.$
[…]
Example 3.2.
Let $(M,g_{M},J)$ be a 4- dimensional Kähler manifold, where $M=\{(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}:x_{i}\neq 0,i=1,...,4\},$ with Riemannian metric
$g_{M}=e^{-2x_{2}}dx_{1}^{2}+dx_{2}^{2}+e^{-2x_{2}}dx_{3}^{2}+dx_{4}^{2}$
and $J(x_{1},x_{2},x_{3},x_{4})=(-x_{2},x_{1},-x_{4},x_{3}).$ Let $(N,g_{N})$ be a 4-dimensional Riemannian manifold, where $N=\{(y_{1},y_{2},y_{3},y_{4})\in\operatorname{\mathbb{R}}^{4}:y_{j}\in\operatorname{\mathbb{R}}\forall j=1,...,4\}$ with Riemannian metric
$g_{N}=dy_{1}^{2}+dy_{2}^{2}+dy_{3}^{2}+dy_{4}^{2}.$
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a map, defined as
$F(x_{1},x_{2},x_{3},x_{4})=(cosx_{2},sinx_{2},x_{4},0).$
Then by direct computation, we get
$kerF_{*}=span\Big{\{}U_{1}=e_{1},U_{2}=e_{3}\Big{\}}$
and
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=e_{2},X_{2}=e_{4}\}.$
Here $\{e_{1}=e^{x_{2}}\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=e^{x_{2}}\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}}\}$ is an orthonormal basis of $T_{p}M$, for any $p\in M.$
Then
$F_{*}(X_{1})=-sinx_{2}e_{1}^{\prime}+cosx_{2}e_{2}^{\prime},$ $F_{*}(X_{2})=e_{3}^{\prime},$
where $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}}\}$ is the standard orthonormal basis of $T_{F(p)}N.$ We can easily check that
$g_{M}(X_{i},X_{i})=g_{N}(F_{*}{X_{i}},F_{*}{X_{i}})$
$\forall i=1,2$ and $J(U_{1})=X_{1},J(U_{2})=X_{2}.$ Thus $F$ is a Langrangian Riemannian map.
To show that $F$ is a Clairaut Langragian Riemannian map, we need to find a smooth function $f$ on $M$ such that $T_{U}U=-g(U,U)\nabla f$ for $U\in\Gamma(kerF_{*}).$
For Riemannian metric $g_{M},$ Christoffel symbols are given by
$\Gamma^{2}_{11}=e^{-2x_{2}},\Gamma^{1}_{12}=-1=\Gamma^{1}_{21},\Gamma^{3}_{23}=-1=\Gamma^{3}_{32},\Gamma^{2}_{33}=e^{-2x_{2}}$ and rest of Christoffel symbols are zero.
The covariant derivatives are calculated as
$\nabla_{U_{1}}U_{1}=X_{1},\nabla_{U_{1}}X_{1}=-U_{1},\nabla_{U_{2}}U_{2}=X_{1},\nabla_{U_{2}}X_{1}=-U_{2}$ and rest are zero.
Gradient of any smooth function $f$ is given by
$\nabla f=\sum_{i,j=1}^{4}g^{ij}\frac{\partial f}{\partial x_{i}}\frac{\partial}{\partial x_{j}}.$
Therefore, we get $\nabla f=-X_{1}.$
Using (3.1), (3.2), (3.3) and above covariant derivatives, we get
$Ric(X_{1},X_{1})=-2,Ric(X_{2},X_{2})=-2,$ and rest are zero.
Since $TM=kerF_{*}\oplus(kerF_{*})^{\perp},$ for any $Y_{i}\in\Gamma(TM),Y_{i}=a_{i}U_{1}+b_{i}U_{2}+c_{i}X_{1}+d_{i}X_{2},i=1,...,4,$ we have
$(L_{Y_{1}}g_{M})(Y_{2},Y_{3})=g_{M}(\nabla_{Y_{2}}Y_{1},Y_{3})+g_{M}(\nabla_{Y_{3}}Y_{1},Y_{2}).$
Substituting the values of covariant derivatives, we obtain
$\begin{split}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})=\frac{1}{2}(a_{1}a_{2}b_{3}+c_{2}c_{1}b_{3}+a_{3}a_{1}b_{2}+c_{3}c_{1}b_{2})-a_{2}b_{1}a_{3}-c_{2}b_{1}c_{3},\end{split}$ (3.26)
$Ric(Y_{2},Y_{3})=-2b_{2}b_{3}-2d_{2}d_{3},$ (3.27)
and
$g_{M}(Y_{2},Y_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}).$ (3.28)
Let $(M,g_{M},J)$ admits a Ricci soliton, then
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0.$
Using (3.26), (3.27) and (3.28) in the above equation, we get
$\lambda=\frac{A_{1}-A_{2}}{2(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{2})},~{}a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}\neq 0,$ (3.29)
where
$A_{1}=2(a_{2}b_{1}a_{3}+c_{2}b_{1}c_{3}+2b_{2}b_{3}+2d_{2}d_{3}),$
$A_{2}=a_{1}a_{2}b_{3}+c_{1}c_{2}b_{3}+a_{1}a_{3}b_{2}+c_{1}c_{3}b_{2},$
[…]
@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$ Then Ricci tensors on $(M,g_{M},J)$ are
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)$ (4.1)
$\begin{split}Ric(F_{*}X,E)=&Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,QE)+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}F_{*}PE}S)_{JF_{*}X}F_{*}X_{j},F_{*}X_{j}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{X_{j}}}S)_{JF_{*}X}PE,F_{*}X_{j}\Big{)}+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{QE}JF_{*}X)\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})JF_{*}X,e_{k}),\end{split}$ (4.2)
$\begin{split}Ric(D,E)=&Ric^{rangeF_{*}}(PD,PE)-\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}\\ &+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PD}}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PD,e_{k})QE,e_{k})+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PE}}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})QD,e_{k})\\ &+Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{M}(\nabla^{M}_{QD}\nabla^{M}g,QE)\Big{)}\end{split}$ (4.3)
@@
Making use of proposition 5.1 [7], we get (4.1) and (4.2). For the proof of (4.3), we proceed as follows.
Since $M$ is a Kähler manifold, we have
$Ric(D,E)=Ric(JD,JE)~{}~{}\forall D,E\in\Gamma(TM).$ (4.4)
@@
$Ric(PD,PE)=\sum_{j=r+1}^{m}g_{M}(R(F_{*}{X_{j}},PD)PE,F_{*}{X_{j}})+\sum_{k=1}^{n_{1}}g_{M}(R(e_{k},PD)PE,e_{k}).$
Using Theorem 3.1([24]) in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{M}(S_{\nabla^{F\perp}_{e_{k}}e_{k}}PD,PE)+\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}S_{e_{k}}PD,PE)\\ &-\sum_{k=1}^{n_{1}}g_{M}(S_{e_{k}}PD,S_{e_{k}}PE)-\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}PD,S_{e_{k}}PE).\end{split}$
Using Theorem 3.1([7]) in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)+\sum_{k=1}^{n_{1}}g_{M}(PD,PE)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{e_{k}}e_{k})\\ &-\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}(e_{k}(g)PD),PE)+\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}PD,e_{k}(g)PE)\\ &-\sum_{k=1}^{n_{1}}(e_{k}(g))^{2}g_{M}(PD,PE).\end{split}$
@@
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{M}(PD,PE)\Big{(}||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}.\end{split}$ (4.6)
@@
$\begin{split}Ric(PD,QE)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*F_{*}PD}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{n_{1}}g_{M}(R^{F\perp}(PD,e_{k})QE,e_{k}),\end{split}$ (4.7)
$\begin{split}Ric(PE,QD)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*F_{*}PE}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{n_{1}}g_{M}(R^{F\perp}(PE,e_{k})QD,e_{k})\end{split}$ (4.8)

### v2

@@ 2.2 Some results on Clairaut anti-invariant Riemannian maps to Kähler manifolds
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is totally geodesic. Then
@@
where $\alpha^{\prime}=\frac{1}{r}.$
@@
Since $\eta$ is horizontal and $(kerF_{*})^{\perp}$ is totally geodesic, we get
@@
$\frac{1}{2}(L_{F_{*}(\nabla f)}g_{N})(F_{*}JU,F_{*}CX)-\alpha^{\prime}Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=0.$
@@
Let $(M,g_{M},J)$ be a 6-dimensional Kähler manifold, where $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}:x_{i}\neq 0,i=1,\ldots,6\},$ with Riemannian metric
$g_{M}=e^{-2x_{4}}dx_{1}^{2}+e^{-2x_{4}}dx_{2}^{2}+e^{-2x_{4}}dx_{3}^{2}+dx_{4}^{2}+dx_{5}^{2}+dx_{6}^{2},$
@@
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=e_{2},X_{2}=e_{4},X_{3}=e_{5},X_{4}=e_{6}\},$
where
$\{e_{1}=e^{x_{4}}\frac{\partial}{\partial x_{1}},e_{2}=e^{x_{4}}\frac{\partial}{\partial x_{2}},e_{3}=e^{x_{4}}\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}},e_{5}=\frac{\partial}{\partial x_{5}},e_{6}=\frac{\partial}{\partial x_{6}}\}$
@@
$F_{*}(X_{1})=e_{2}^{\prime},$ $F_{*}(X_{2})=e_{4}^{\prime},F_{*}(X_{3})=e_{1}^{\prime},F_{*}(X_{4})=e_{6}^{\prime},$
where $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=\frac{\partial}{\partial y_{6}}\}$ represents the standard orthonormal basis of $T_{F(p)}N.$
Also, it can be easily verified that for $i=1,...,4$
@@
and $J(U_{1})=X_{1},J(U_{2})=X_{2},J(X_{1})=-U_{1},J(X_{2})=-U_{2},J(X_{3})=X_{4},J(X_{4})=-X_{3}.$ Thus $F$ is an anti-invariant Riemannian map.
@@
$\Gamma^{1}_{14}=-1=\Gamma^{1}_{41},\Gamma^{4}_{11}=e^{-2x_{4}},\Gamma^{4}_{22}=e^{-2x_{4}},\Gamma^{4}_{33}=e^{-2x_{4}},\Gamma^{2}_{24}=-1=\Gamma^{2}_{42},\Gamma^{3}_{34}=-1=\Gamma^{3}_{43}$ and rest of them are zero.
@@
$\nabla_{U_{1}}U_{1}=X_{2},\nabla_{U_{1}}X_{2}=-U_{1},\nabla_{X_{1}}X_{1}=X_{2},\nabla_{U_{2}}U_{2}=X_{2},\nabla_{X_{1}}X_{2}=-X_{1},\nabla_{U_{2}}X_{2}=-U_{2},$ rest of them are zero.
Ricci curvatures are obtained as $Ric(U_{1},U_{1})=3,Ric(U_{2},U_{2})=1,Ric(U_{1},U_{2})=-1,Ric(X_{1},X_{1})=-2,Ric(X_{2},X_{2})=-2$ remaining Ricci tensors are zero.
@@
$\begin{split}(L_{Z_{1}}g_{M})(Z_{2},Z_{3})=&\frac{1}{2}(a_{1}a_{2}d_{3}+a_{1}a_{3}d_{2}+c_{3}c_{1}d_{2}+b_{3}b_{1}d_{2}+c_{2}c_{1}d_{3}+b_{2}b_{1}d_{3})\\ &-a_{2}a_{3}d_{1}-c_{3}c_{2}d_{1}-b_{2}b_{3}d_{1},\end{split}$ (3.21)
$Ric(Z_{2},Z_{3})=3a_{2}a_{3}+b_{2}b_{3}-a_{2}b_{3}-2c_{2}c_{3}-2d_{2}d_{3},$ (3.22)
@@
$A_{1}=2(3a_{2}a_{3}+b_{2}b_{3}-a_{2}b_{3}-2c_{2}c_{3}-2d_{2}d_{3}),$
$A_{2}=(a_{1}a_{2}d_{3}+a_{1}a_{3}d_{2}+c_{3}c_{1}d_{2}+b_{3}b_{1}d_{2}+c_{2}c_{1}d_{3}+b_{2}b_{1}d_{3})-2a_{2}a_{3}d_{1}-2c_{3}c_{2}d_{1}-2b_{2}b_{3}d_{1}.$
For some particular values of $a_{i}^{\prime}s,b_{i}^{\prime}s,c_{i}^{\prime}s,d_{i}^{\prime}s,l_{i}^{\prime}s,\text{and}~{}h_{i}^{\prime}s,$ the behaviour of Ricci soliton will be shrinking, steady, and expanding accorrding to $\lambda<0,\lambda=0,~{}\text{and}~{}\lambda>0$ respectively.
@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$ Then Ricci tensors on $(N,g_{N},J^{\prime})$ are
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)+(m-r)g_{N}(\nabla^{N}g,\nabla^{F\perp}_{J^{\prime}F_{*}X}J^{\prime}F_{*}Y)$ (4.1)
$\begin{split}Ric(F_{*}X,E)=&Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,QE)+\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}F_{*}PE}S)_{J^{\prime}F_{*}X}F_{*}X_{j},F_{*}X_{j}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{X_{j}}}S)_{J^{\prime}F_{*}X}PE,F_{*}X_{j}\Big{)}+(m-r)g_{N}(\nabla^{N}g,\nabla^{F\perp}_{QE}J^{\prime}F_{*}X)\\ &-\sum_{k=1}^{r_{1}}g_{N}(R^{F\perp}(PE,e_{k})J^{\prime}F_{*}X,e_{k}),\end{split}$ (4.2)
$\begin{split}Ric(D,E)=&Ric^{rangeF_{*}}(PD,PE)-\sum_{k=1}^{r_{1}}g_{N}(PD,PE)\Big{(}||\nabla^{N}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}\\ &+\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PD}}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{r_{1}}g_{N}(R^{F\perp}(PD,e_{k})QE,e_{k})+\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PE}}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{r_{1}}g_{N}(R^{F\perp}(PE,e_{k})QD,e_{k})\\ &+Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{N}(\nabla^{N}_{QD}\nabla^{N}g,QE)\Big{)}\end{split}$ (4.3)
@@
Making use of Proposition 5.1 of [11], we get (4.1) and (4.2). For the proof of (4.3), we proceed as follows.
Since $N$ is a Kähler manifold, we have
$Ric(D,E)=Ric(JD,JE)~{}~{}\forall D,E\in\Gamma(rangeF_{*})^{\perp}.$ (4.4)
@@
$Ric(PD,PE)=\sum_{j=r+1}^{m}g_{N}(R(F_{*}{X_{j}},PD)PE,F_{*}{X_{j}})+\sum_{k=1}^{n_{1}}g_{N}(R(e_{k},PD)PE,e_{k}).$
Using Theorem 3.1 of [32] in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{N}(S_{\nabla^{F\perp}_{e_{k}}e_{k}}PD,PE)+\sum_{k=1}^{n_{1}}g_{N}(\nabla^{N}_{e_{k}}S_{e_{k}}PD,PE)\\ &-\sum_{k=1}^{n_{1}}g_{N}(S_{e_{k}}PD,S_{e_{k}}PE)-\sum_{k=1}^{n_{1}}g_{N}(\nabla^{N}_{e_{k}}PD,S_{e_{k}}PE).\end{split}$
Using Theorem 3.1 of [11] in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)+\sum_{k=1}^{n_{1}}g_{N}(PD,PE)g_{N}(\nabla^{N}g,\nabla^{F\perp}_{e_{k}}e_{k})\\ &-\sum_{k=1}^{n_{1}}g_{N}(\nabla^{N}_{e_{k}}(e_{k}(g)PD),PE)+\sum_{k=1}^{n_{1}}g_{N}(\nabla^{N}_{e_{k}}PD,e_{k}(g)PE)\\ &-\sum_{k=1}^{n_{1}}(e_{k}(g))^{2}g_{N}(PD,PE).\end{split}$
@@
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{N}(PD,PE)\Big{(}||\nabla^{N}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}.\end{split}$ (4.6)
@@
$\begin{split}Ric(PD,QE)=&\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}F_{*}PD}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{n_{1}}g_{N}(R^{F\perp}(PD,e_{k})QE,e_{k}),\end{split}$ (4.7)
$\begin{split}Ric(PE,QD)=&\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}F_{*}PE}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{n_{1}}g_{N}(R^{F\perp}(PE,e_{k})QD,e_{k})\end{split}$ (4.8)